# Emergence in graphs with near-extreme constraints

### v2

@@ 3. Proof of Theorem 1
We will prove a slightly more quantitative version of Theorem 1:
@@ 3.1. Strategy
The proofs of Theorems 6, 7 and 9 all follow the same general outline. We will present the proof of Theorem 6 in full detail. The subsequent proofs of Theorems 7 and 9 will be somewhat abbreviated, concentrating on what is different in those cases.
Using the fact that the unique entropy maximizing graphon $g_{0}$ at $(e,0)$ is symmetric bipodal, we show that the optimizing graphons at points near the boundary have the same general structure away from a small exceptional set. Specifically, we partition the unit interval into subsets $I_{1}$, $I_{2}$ and $I_{3}$ such that the columns $g_{x}$ of the optimal graphon are $L^{2}$-close to the columns of the first pode of $g_{0}$ when $x\in I_{1}$ and are $L^{2}$-close to the columns of the second pode of $g_{0}$ when $x\in I_{2}$, and where $I_{3}$ has small measure. At this stage, we do not have any control over $g_{x}$ when $x\in I_{3}$.
@@
[…]
Using the pointwise equations (9), we bound the maximum fluctuation of $g(x,y)$ in each rectangle $I_{i}\times I_{j}$ by a multiple of the maximum fluctuation in a neighboring rectangle. Combining these results, the maximum fluctuation in each rectangle is bounded by a small multiple of itself, and so must be zero. That is, our optimal graphon must be exactly bipodal.
The space of bipodal graphons with given values of $(e,t)$ is only 2-dimensional. Using ordinary calculus, we determine that the entropy is maximized when the graphon is symmetric.
@@ 3.2. Defining approximate podes
To see this, we invoke the compactness of the space of reduced graphons in the cut metric. A subsequence must converge to a limit $g_{\infty}$ in the cut metric. By the semicontinuity of the entropy and the fact that the limit of the entropies of the entropy maximizers is at least $S(g_{0})$, we must have $S(g_{\infty})\geq S(g_{0})$. But $g_{0}$ is the unique entropy maximizer at $(e,0)$, so $g_{\infty}=g_{0}$.
The entire sequence $\{g_{i}\}$, and not just a subsequence, must converge to $g_{0}$. If it did not, we could find a subsequence where all points were bounded away from $g_{0}$ in the cut metric. Applying the previous argument to this subsequence would then yield a contradiction.
The entropy $S(g)$ is continuous in $L^{2}$ but only semicontinuous in the cut metric. Limits in the cut metric that are not limits in $L^{2}$ are associated with jumps in the entropy. Since $S(g_{0})=\lim S(g_{i})$, the convergence of $g_{i}$ to $g_{0}$ must occur in $L^{2}$ as well as in the cut metric. That is, for every $\epsilon>0$ there is a $\delta>0$ such that, for all $t<\delta$ and all optimal graphons $g$ at $(e,t)$, $\|g-g_{0}\|_{L^{2}}<\epsilon$. (Note that we have not assumed that the optimal graphon $g$ is unique. That will be proven in due course.)
@@
[…]
except on a set of $x$’s of measure $\epsilon$ or less. Call that exceptional set $I_{3}$. Let $I_{1}$ and $I_{2}$ be the intersection of $I_{3}^{c}$ with $[0,1/2]$ and $[1/2,1]$, respectively. That is, we have broken the unit interval into three pieces $I_{1}$, $I_{2}$, $I_{3}$, such that:
  • For all $x\in I_{1}$, $\|g_{x}-C_{1}\|<\sqrt{\epsilon}$.
  • For all $x\in I_{2}$, $\|g_{x}-C_{2}\|<\sqrt{\epsilon}$.
@@
Here $C_{1}$ and $C_{2}$ are the columns of $g_{0}$ on the two podes, namely $2e$ times the indicator functions of $[1/2,1]$ and $[0,1/2]$, respectively. We will refer to the sets $I_{1}$, $I_{2}$ and $I_{3}$ as podes, even though we are not assuming that the graphon $g$ is exactly tripodal.
@@ 3.3. Variational equations
We examine the terms in our variational equations (9). The first two derivatives of the function $H$ are
@@ 3.4. Maximizing worth and eliminating $I_{3}$
[…]
Up to small deviations due to the differences between $g$ and $g_{0}$, maximizing $W(C)$ amounts to taking $C(y)$ to be constant on $[0,1/2]$ and constant on $[1/2,1]$ and choosing $a$ and $b$ to maximize
@@
If $a$ is substantially nonzero (say, bigger than $1/\sqrt{\beta}$), then $e\beta a$ is gigantic and $b$ is extremely close to zero, being $O(\exp(-\sqrt{\beta}))$. This makes $e\beta b$ tiny so $H^{\prime}(a)\approx\alpha\approx H^{\prime}(2e)$ and $a\approx 2e$. Similarly, if $b$ is substantially nonzero then $a$ is tiny and $b\approx 2e$. In both those cases, $W(C)\approx\frac{1}{2}H(2e)-eH^{\prime}(2e)=-\frac{1}{2}\ln(1-2e)$. The third possibility is that $a$ and $b$ are both tiny, but in that case $W(C)\approx 0$, which is strictly less than $-\frac{1}{2}\ln(1-2e)$.
The upshot is that there are three stationary points of (20) but only two maxima, one that resembles $g_{x}$ for $x\in I_{1}$ and one that resembles $g_{x}$ for $x\in I_{2}$. Since every column $g_{x}$ with $x\in I_{3}$ must be a worth-maximizer, and since every worth-maximizer must come close to maximizing (20), every column $g_{x}$ with $x\in I_{3}$ resembles $g_{x}$ for $x\in I_{1}$ or $I_{2}$. We can then reassign the points of $I_{3}$ to $I_{1}$ or $I_{2}$ depending on the nature of $g_{x}$.
@@ 3.5. Exact bipodality
Our next step is to upgrade our $L^{2}$ estimates on the forms of the different columns into pointwise estimates. Thanks to each column of $g$ being $L^{2}$-close to a column of $g_{0}$, the function $G(x,y)=\langle g_{x}|g_{y}\rangle$ is close to $2e^{2}$ on $I_{1}\times I_{1}$ and on $I_{2}\times I_{2}$. By (9), this forces $g(x,y)$ to be exponentially small (in $\beta$) in these quadrants. This in turn makes $G(x,y)$ exponentially small on $I_{1}\times I_{2}$ and $I_{2}\times I_{1}$, which means that $H^{\prime}(g)$ is exponentially close to $\alpha$ in these quadrants, and therefore that $g(x,y)$ is pointwise close to constant in these rectangles.
@@
(25) $-H^{\prime\prime}(A_{0})\Delta_{A}\leq 4Ac\beta\Delta_{A}+4D(1-c)\beta\Delta_{D},$
@@
(30) $-H^{\prime\prime}(D_{0})\Delta D\leq 2\beta(cA+(1-c)B)\Delta_{D}+2\beta D(c\Delta_{A}+(1-c)\Delta_{B}).$
@@ 3.6. Symmetric bipodality
All that remains is showing that the best bipodal graphon is symmetric, with pode sizes $\frac{1}{2}$ and $\frac{1}{2}$ and with $A=B$. This requires extensive calculations but no sophisticated analysis. Ultimately, it is just a (grungy) problem in multivariable calculus.
@@
Now consider the effect of having $t$ nonzero. Having the graphon nonzero on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ provides additional entropy of order $t\ln(1/t)$. Shifting the value of the graphon on $I_{1}\times I_{2}$ by an $O(t)$ amount changes the entropy by an additional $O(t)$, but since this is small compared to $t\ln(1/t)$, $S(e,t,c)-S(e,0,c)$ is still $O(t\ln(1/t))$. In order to overcome the $-2\ln(1-2e)\Delta c^{2}$ cost, we must have $\Delta c=O(\sqrt{t\ln(1/t)})$. Since $t\sim\exp(-2e^{2}\beta)$, this means that $\Delta c$ must be exponentially small in $\beta$ and in particular that $\beta\Delta c$ is a small parameter.
@@
since $\beta=O(\ln(1/t))$. This possible entropy gain from having $\Delta c\neq 0$ is much smaller than the $-2\ln(1-2e)\Delta c^{2}$ cost, so the optimal value of $\Delta c$ is exactly zero. That is, we must have $c=1/2$.
@@
where we have used the fact that $D=2e-A$. Since $A$ was exponentially small in $\beta$, $\beta$ must scale as $\ln(1/t)$. The entropy is
@@
The Boltzmann entropy $\mathbb{B}(e,t)$ is equal to the Shannon entropy $S(g)$ of the optimal graphon at $(e,t)$, so $\Delta\mathbb{B}:=\mathbb{B}(e,t)-\mathbb{B}(e,0)=S(g)-\frac{1}{2}H(2e)$. Since $A\approx(t/3e^{2})$ and $A\approx\exp(-2e^{2}\beta)$, $\beta\approx\frac{-1}{2e^{2}}\ln(t/3e^{2})\sim\ln(1/t)$. ∎
@@ 4.1. Defining approximate podes
As we approach the scallop, any sequence of entropy maximizers must converge to $g_{0}$ in $L^{2}$ by exactly the same argument as in the proof of Theorem 6.
As before, we pick a sufficiently small value of $\epsilon$ and consider values of $t$ small enough that $\|g-g_{0}\|_{L^{2}}<\epsilon$ for each optimal graphon $g$. Let $I_{1}$, $I_{2}$, $\ldots$, $I_{n+2}$ be the subsets of the podes of $g_{0}$ for which $g_{x}$ lies within $\sqrt{\epsilon}$ in $L^{2}$ of the corresponding column of $g_{0}$ and let $I_{n+3}$ be the exceptional set where $g_{x}$ is not close to the corresponding column of $g_{0}$. Note that $I_{n+3}$ may contain points $x$ where $g_{x}$ is close to a different column of $g_{0}$. Those points will soon be reassigned.

### v3

@@ 3. Proof of Theorem 1
We now prove a slightly more quantitative version of Theorem 1:
@@ 3.1. Strategy
The proofs of Theorems 14, 17 and 20 all follow the same general outline. We will present the proof of Theorem 14 in full detail. The subsequent proofs of Theorems 17 and 20 will be somewhat abbreviated, concentrating on what is different in those cases.
Using the fact that the unique entropy maximizing graphon $g_{0}$ at $(e,0)$ is symmetric bipodal, we show that the optimizing graphons at points near the boundary have the same general structure away from an exceptional set of small area. Specifically, we partition the unit interval into subsets $I_{1}$, $I_{2}$ and $I_{3}$ such that the columns $g_{x}$ of the optimal graphon are $L^{2}$-close to the columns of the first pode of $g_{0}$ when $x\in I_{1}$ and are $L^{2}$-close to the columns of the second pode of $g_{0}$ when $x\in I_{2}$, and where $I_{3}$ has small measure. At this stage, we do not have any control over $g_{x}$ when $x\in I_{3}$.
@@
We then study the worth functional $W(C)$. The dependence of this functional on the graphon $g$ comes via the integral $\iint C(y)C(z)g(y,z)\,dy\,dz$. Since $C$ is bounded, and since we know $g(y,z)$ away from a set of small measure, we have good control over $W(C)$. We determine that a worth-maximizing column can only take one of two approximate forms, namely those exhibited by $g_{x}$ for $x\in I_{1}$ and for $x\in I_{2}$. We then reassign each point $x\in I_{3}$ to $I_{1}$ or $I_{2}$, depending on which worth-maximizing form $g_{x}$ takes. The result is then a graphon with two (approximate) podes.
Using the pointwise equations (16), we bound the variation of $g(x,y)$ in each rectangle $I_{i}\times I_{j}$ by a multiple of the variation in a neighboring rectangle. Combining these results, the variation in each rectangle is bounded by a small multiple of itself, and so must be zero. That is, our optimal graphon must be exactly bipodal.
The space of bipodal graphons with given values of $(e,t)$ is only 2-dimensional. Using ordinary 2-dimensional calculus, we determine that the entropy $S(g)$ is maximized when the graphon is symmetric.
To account for the fact that the Lagrange multipliers $\alpha$ and $\beta$ are only defined for almost every $t$ and not necessarily for every $t$, we analyze optimal graphons in two passes. In the first pass we use the worth function and pointwise equations (16), as outlined above, to establish that the optimal graphon is symmetric bipodal for almost every $t$ that is sufficiently small. This is the bulk of the proof of Theorem 14.
This shows that the Boltzmann entropy is almost everywhere equal to the Shannon entropy of a symmetric bipodal graphon, which is a differentiable function of $t$ for each $e$. By Theorem 13, this then implies that there are no singular values of $t$ and that our arguments apply at all sufficiently small values of $t$.
The same two-pass argument applies to the proofs of Theorems 17 and 20, only replacing “symmetric bipodal” with the particular graphon symmetry described in those theorems. Specifically, in the first pass we show that the Boltzmann entropy is almost everywhere equal to the maximum Shannon entropy among multipodal graphons of a certain sort. The solution to the resulting finite-dimensional maximization problem yields a smooth function of $t$. Theorem 13 then says that our results apply at every $t$.
Remark 15.
As is standard when working with $L^{2}$ spaces, the proofs of Theorems 14, 17 and 20 are written as if our graphons were actual functions on $0,1]^{2}$. But in fact they are equivalence classes of functions that agree away from a set of measure zero. This has several consequences, none of which materially affect the flow of the proofs:
  • When we use the variational equations (16), the results apply almost everywhere, not literally everywhere. Whenever we use those equations to compute an upper or lower bound on a graphon, that bound should always be understood to mean “apart from on a set of measure zero”.
  • When we speak of the “maximum” value of a graphon $g$ on a region, we actually mean the essential supremum of the function, namely the smallest number $M$ such that $g(x,y)\leq M$ on a set of full measure. The “minimum” is similar.
  • Since we are working with functions mod sets of measure zero, we are free to change the value of our graphon on sets of measure zero whenever we wish. In this way, we could make the variational equations apply everywhere, or we could make the maximum of a function equal the essential supremum. We could, but we won’t actually subject the reader to such painstaking bookkeeping!
Instead, we will not keep track of sets of measure zero in these proofs, such as deciding whether a pode contains its endpoints. All the important properties of graphons (or at least all the properties considered in this paper) are based on integrals, for which sets of measure zero don’t matter at all.
Having explained the process in all three settings, we return our focus to the first pass.
@@ 3.2. Defining approximate podes
To see this, we invoke the compactness of the space of reduced graphons in the cut metric. A subsequence must converge to a limit $g_{\infty}]$ in the cut metric. Lemma 2.1 in [14] proves that $S$ is upper-semicontinuous on $\widetilde{\mathcal{W}}$ in the cut metric $\delta_{cut}$. This implies that the limit of the entropies of the entropy maximizers is at least $S([g_{0}])$, so we must have $S([g_{\infty}])\geq S([g_{0}])$. But $g_{0}]$ is the unique entropy maximizing reduced graphon at $(e,0)$, so $g_{\infty}]=[g_{0}]$.
The entire sequence $\{[g_{i}]\}$, and not just a subsequence, must converge to $g_{0}]$. If it did not, we could find a subsequence where all points were bounded away from $g_{0}]$ in the cut metric $\delta_{cut}$. Applying the previous argument to this subsequence would then yield a contradiction. That is, after applying appropriate measure-preserving transformations of $[0,1]$, the sequence $\{g_{i}\}$ of entropy maximizing graphons must converge to $g_{0}$ in the cut distance $d_{cut}$.
Let $I_{1}=[0,1/2]$ and $I_{2}=[1/2,1]$ be the two podes of $g_{0}$. By the definition of the cut distance, the average value of $g_{i}$ must converge to 0 on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ and to $2e$ on $I_{1}\times I_{2}$ and $I_{2}\times I_{1}$. The variance of $g_{i}$ must go to zero on each of these rectangles, or else $\lim S(g_{i})$ would be strictly less than $S(g_{0})$. Having $\lim S(g_{i})<S(g_{0})$ is impossible because there exist explicit symmetric bipodal graphons with $t\to 0$ whose entropies give a lower bound for $S(g_{i})$ and whose entropies approach $S(g_{0})$ as $t\to 0$.
Since the mean of $g_{i}$ in each rectangle approaches the value of $g_{0}$ and since the variance goes to zero, $\{g_{i}\}$ approaches $g_{0}$ in $L^{2}$. That is, for every $\epsilon>0$ there is a $\delta>0$ such that, for all $t<\delta$ and all optimal graphons $g$ at $(e,t)$, applying a measure-preserving transformation of $[0,1]$ to $g$ we have $\|g-g_{0}\|_{2}<\epsilon$. (Note that we have not assumed that the optimal graphon $g$ is unique. That will be proven in due course.)
@@
(31) $\int_{0}^{1}\big{(}g(x,y)-g_{0}(x,y)\big{)}^{2}dy<\epsilon,$
except on a set of $x$’s of measure $\epsilon$ or less. Call that exceptional set $I_{3}$. Let $I_{1}$ and $I_{2}$ be the intersection of $I_{3}^{c}$ with $[0,1/2]$ and $[1/2,1]$, respectively. Let $C_{1}$ and $C_{2}$ to be the columns of $g_{0}$ on the two podes, namely $2e$ times the indicator functions of $[1/2,1]$ and $[0,1/2]$, respectively. We have broken the unit interval into three pieces $I_{1}$, $I_{2}$, $I_{3}$, such that:
  • For all $x\in I_{1}$, $\|g_{x}-C_{1}\|_{2}<\sqrt{\epsilon}$.
  • For all $x\in I_{2}$, $\|g_{x}-C_{2}\|_{2}<\sqrt{\epsilon}$.
@@
We will refer to the sets $I_{1}$, $I_{2}$ and $I_{3}$ as podes, even though we are not assuming that the graphon $g$ is exactly tripodal.
@@ 3.3. Variational equations
We now use the pointwise variation equations (16) to get some preliminary estimates on $g(x,y)$. The first two derivatives of the function $H$ are
@@ 3.4. Maximizing worth and eliminating $I_{3}$
  • The entropy term $2\int_{0}^{1}H(C(y))\,dy$ is bounded above by $H(a)+H(b)$, thanks to $H^{\prime\prime}$ being everywhere negative.
  • The term $2\alpha\int_{0}^{1}C(y)\,dy$ is exactly $\alpha(a+b)$.
  • The term $\beta\iint C(y)C(z)g(y,z)\,dy\,dz$ is approximately $e\beta ab$.
Recall that the worth of a column is
(37) $W(C)=2\int_{0}^{1}H(C(y))\,dy-2\alpha\int_{0}^{1}C(y)\,dy-\beta\iint C(y)C(z)g(y,z)\,dy\,dz.$
If $g$ were equal to $g_{0}$, maximizing $W(C)$ would involve taking $C(y)$ to be constant on $[0,1/2]$ and constant on $[1/2,1]$ and choosing $a$ and $b$ to maximize
@@
(Because of the small differences between $g$ and $g_{0}$, this procedure only gives approximate worth maximizers, but that is enough for our purposes.)
@@
If $a$ is substantially nonzero (say, bigger than $1/\sqrt{\beta}$), then $e\beta a$ is gigantic and $b$ is extremely close to zero, being $O(\exp(-\sqrt{\beta}))$. This makes $e\beta b$ tiny so $H^{\prime}(a)\approx\alpha\approx H^{\prime}(2e)$ and $a\approx 2e$. Similarly, if $b$ is substantially nonzero then $a$ is tiny and $b\approx 2e$. In both those cases, $W(C)\approx H(2e)-2eH^{\prime}(2e)=-\ln(1-2e)$. The third possibility is that $a$ and $b$ are both tiny, but in that case $W(C)\approx 0$, which is strictly less than $\ln(1-2e)$.
The upshot is that there are three stationary points of (38) but only two maxima, one that resembles $g_{x}$ for $x\in I_{1}$ and one that resembles $g_{x}$ for $x\in I_{2}$. Since every column $g_{x}$ with $x\in I_{3}$ must be a worth-maximizer, and since every worth-maximizer must come close to maximizing (38), every column $g_{x}$ with $x\in I_{3}$ is close in $L^{2}$ to the columns for $x\in I_{1}$ or $I_{2}$. We can then reassign the points of $I_{3}$ to $I_{1}$ or $I_{2}$ depending on the nature of $g_{x}$.
@@ 3.5. Exact bipodality
Our next step is to upgrade our $L^{2}$ estimates on the forms of the different columns into pointwise estimates. Thanks to each column of $g$ being $L^{2}$-close to a column of $g_{0}$, the function $G(x,y)=\langle g_{x}|g_{y}\rangle$ is pointwise close to $2e^{2}$ on $I_{1}\times I_{1}$ and on $I_{2}\times I_{2}$. By (16), this forces $g(x,y)$ to be exponentially small (specifically, $\exp(-\Theta(\beta)))$ in these quadrants. This in turn makes $G(x,y)$ exponentially small on $I_{1}\times I_{2}$ and $I_{2}\times I_{1}$, which means that $H^{\prime}(g)$ is exponentially close to $\alpha$ in these quadrants, and therefore that $g(x,y)$ is pointwise close to constant in these rectangles.
@@
(43) $H^{\prime\prime}(A_{0})\Delta_{A}\leq 4Ac\beta\Delta_{A}+4D(1-c)\beta\Delta_{D},$
@@
(48) $H^{\prime\prime}(D_{0})\Delta D\leq 2\beta(cA+(1-c)B)\Delta_{D}+2\beta D(c\Delta_{A}+(1-c)\Delta_{B}).$
@@ 3.6. Symmetric bipodality
All that remains is showing that the best bipodal graphon is symmetric, with pode sizes $\frac{1}{2}$ and $\frac{1}{2}$ and with $A=B$. This requires extensive calculations but no sophisticated analysis. Ultimately, it is just a (grungy) problem in multivariable calculus as follows.
@@
Now consider the effect of having $t$ nonzero. Having the graphon nonzero on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ provides additional entropy of order $t\ln(1/t)$. Shifting the value of the graphon on $I_{1}\times I_{2}$ by an $O(t)$ amount changes the entropy by an additional $O(t)$, but since this is small compared to $t\ln(1/t)$, S(e,t,c)S(e,0,c)S(e,t,c)-S(e,0,c) is still $O(t\ln(1/t))$. In order to overcome the $2\ln(1-2e)\Delta c^{2}$ cost, we must have $\Delta c=O(\sqrt{t\ln(1/t)})$. Since $t\sim\exp(-2e^{2}\beta)$, this means that $\Delta c$ must be exponentially small in $\beta$ and in particular that $\beta\Delta c$ is a small parameter.
@@
since $\beta=O(\ln(1/t))$. This possible entropy gain from having $\Delta c\neq 0$ is much smaller than the $2\ln(1-2e)\Delta c^{2}$ cost, so the optimal value of $\Delta c$ is exactly zero. That is, we must have $c=1/2$.
@@
where we have used the fact that D=2eAD=2e-A. Since $A$ was exponentially small in $\beta$, $\beta$ must scale as $\ln(1/t)$. The entropy is
@@
The Boltzmann entropy $\mathbb{B}(e,t)$ is equal to the Shannon entropy $S(g)$ of the optimal graphon at $(e,t)$, so $\Delta\mathbb{B}:=\mathbb{B}(e,t)-\mathbb{B}(e,0)=S(g)-\frac{1}{2}H(2e)$. Since $A\approx(t/3e^{2})$ and $A\approx\exp(-2e^{2}\beta)$, $\beta\approx\frac{-1}{2e^{2}}\ln(t/3e^{2})\sim\ln(1/t)$. This completes the proof of Theorem 14. ∎
Note that we have actually proved something slightly stronger than Theorem 14. We started with an optimal graphon, applied measure preserving transformations of $[0,1]$, and wound up with a symmetric bipodal graphon. That is, the optimal graphon in $\mathcal{W}_{e,t}$ is unique up to group equivalence. However, a reduced graphon in $\widetilde{\mathcal{W}}_{e,t}$ is a weak equivalence class, with all representatives of this class being constrained entropy maximizers. We thus conclude that
Corollary 16.
For all $e<1/2$ and for all $t$ small enough that Theorem 14 applies, any graphon $g\in\mathcal{W}_{e,t}$ that is weakly equivalent to a symmetric bipodal graphon is group equivalent to a symmetric bipodal graphon.
@@ 4. Proof of Theorem 2
  6. (6)
    On the first pass, steps (1–5) only apply at values of $t$ for which the Lagrange multipliers $\alpha$ and $\beta$ are well-defined and finite. Extending the results to all sufficiently small values of $\Delta t$ then follows from Theorem 13.
@@ 4.1. Defining approximate podes
As we approach the scallop, any sequence of entropy maximizers must converge (after applying measure preserving transformations of $[0,1]$) to $g_{0}$ in $L^{2}$ by exactly the same argument as in the proof of Theorem 14.
As before, any sequence of entropy maximizing reduced graphons must converge to $g_{0}]$ in the cut metric, implying than any sequence of entropy maximizing graphons must are equivalent to graphons $g_{i}$ that converge to $g_{0}$ in the cut distance. By the definition of the cut distance, the average value of $g_{i}$ on each rectangle defined by the podes of $g_{0}$ must approach the (constant) value of $g_{0}$ on that rectangle. Since $\lim S(g_{i})=S(g_{0})$, the variance of $g_{i}$ must go to zero on each of these rectangles, so the graphons $g_{i}$ converge to $g_{0}$ in $L^{2}$.
We pick a sufficiently small value of $\epsilon$ and consider values of $t$ small enough that $\|g-g_{0}\|_{L^{2}}<\epsilon$ for each optimal graphon $g$. Let $I_{1}$, $I_{2}$, $\ldots$, $I_{n+2}$ be the subsets of the podes of $g_{0}$ for which $g_{x}$ lies within $\sqrt{\epsilon}$ in $L^{2}$ of the corresponding column of $g_{0}$ and let $I_{n+3}$ be the exceptional set where $g_{x}$ is not close to the corresponding column of $g_{0}$. Note that $I_{n+3}$ may contain points $x$ where $g_{x}$ is close to a different column of $g_{0}$. Those points will soon be reassigned.